# Can LLM-Simulated Practice and Feedback Upskill Human Counselors? A Randomized Study with 90+ Novice Counselors

### v1

@@
Abstract
Training more counselors, from clinical students to peer supporters, can help meet the demand for accessible mental health support; however, current training approaches remain resource-intensive and difficult to scale effectively. Large Language Models (LLMs) offer promising solutions for growing counseling skills training through simulated practice and automated feedback. Despite successes in aligning LLMs with expert-counselor annotations, we do not know whether LLM-based counseling training tools—such as AI patients that simulate real-world challenges and generative AI feedback with suggested alternatives and rationales—actually lead to improvements in novice counselor skill development. We develop CARE, an LLM-simulated practice and feedback system, and randomize 94 novice counselors to practice using an AI patient, either alone or with AI feedback, measuring changes in their behavioral performance, self-assessments, and qualitative learning takeaways. Our results show the practice-and-feedback group improved in their use of reflections and questions (d=0.32-0.39, p$<$0.05). In contrast, the group that practiced with an AI patient alone did not show improvements, and in the case of empathy, actually had worse uses across time (d=--0.52, p=0.001) and when compared against the practice-and-feedback group (d=0.72, p=0.001). Participants’ qualitative self-reflections revealed key differences: the practice-and-feedback group adopted a client-centered approach involving listening to and validating feelings, while the practice-alone group remained solution-oriented but delayed offering suggestions until gathering more information. Overall, these results suggest that LLM-based training systems can promote effective skill development, but that combining both simulated practice and structured feedback is critical.
@@ Abstract
In 2023, 22.8% of U.S. adults (approximately 58.7 million people) experienced a mental illness [1]. Unfortunately, the ability to provide people with effective mental health care is severely limited by shortages of high-quality mental health providers—from psychotherapists and counselors to social workers and peer supporters [2, 3]. While there is increasing interest in exploring direct-to-patient AI systems for some circumstances, and some promising results [4], we expect the demand for human therapists and mental health support to continue to far exceed demand. The limited supply of effective therapy providers is due—at least in part—to the reliance on resource-intensive methods to train helping skills [5] and evidence-based interventions [6, 7]. Traditional methods require access to trainers who can simulate a client interaction [8, 9] and provide expert supervision [10, 11], creating significant barriers to growing high-quality training [12, 13].
@@
Fortunately, AI systems have been increasingly applied to counselor training as a potential solution to these scaling challenges. Recent advances in large language models (LLMs) have enabled the development of LLM-based systems for simulating patients seeking mental health support [14, 15], offering rich opportunities for practice. Simultaneously, AI feedback systems have progressed in automatically assessing counselor behaviors such as empathy, reflections, and active listening [16, 17, 18, 19] and generating suggested responses [20, 21, 22] and explanations [23, 22]. These AI feedback systems target skills from client-centered approaches [24, 25]—empathy, reflections, questions, and active listening—which are relevant to all mental health providers. Meta-analyses have shown these skills strengthen common factors like therapeutic alliance, which are powerful predictors of therapy outcomes, regardless of the eventual evidence-based intervention or therapy modality implemented [26, 27, 28]. In addition to technical advances, prior work evaluating AI feedback systems with novice counselors have tested their effectiveness when used as real-time co-pilots rather than as training systems [29, 21], or evaluated pre-LLM training systems which had limited practice realism and simpler, non-generative feedback mechanisms [30]. Therefore, we have limited evidence explaining how modern, LLM-based training systems—which offer practice simulations resembling challenging real-world behaviors and generative feedback with suggested alternatives and rationales—promote counselor skill development [31, 32]–a holistic concept that includes appropriate use of skill behaviors, accurate self-awareness of one’s capabilities and limitations [33], and an ability to reflect on and learn from experiential practice [34].
To address this gap, we develop CARE, an LLM-based training system for novice counselors, and conduct a randomized experiment to investigate how two different modes of simulated training impact counselor skill development. CARE enables (1) realistic practice with LLM-simulated patients prompted by expert counselors to resemble challenging behaviors [15], and (2) structured feedback from a fine-tuned LLM that identifies strengths and areas for improvement across core counseling skills (e.g., empathy, reflections, questions, suggestions), while also providing explanatory rationale and alternative responses [22]. Importantly, CARE’s feedback evaluates not only when a counseling skill is used but also how well it is used, distinguishing between strong implementations (e.g., reflections that accurately capture client emotions) and implementations needing improvement (e.g., reflections that miss core client concerns). These feedback evaluations are based on established counseling frameworks [5, 35] and expert-annotated examples.

### v2

@@
Abstract.
The growing demand for accessible mental health support requires training more counselors, yet existing approaches remain resource-intensive and difficult to scale. LLMs can realistically simulate patients and generate actionable feedback for training, but their actual impact on novice counselor skill development remains unknown. We developed an LLM-simulated practice and feedback system and conducted a randomized study with 94 novice counselors, comparing practice alone versus practice with feedback. We evaluated behavioral performance, self-efficacy, and qualitative reflections. Results showed the practice-and-feedback group improved in client-centered microskills (reflections, questions), while the practice-alone group showed no improvements. For empathy, the practice-alone group declined over time and performed significantly worse than the feedback group. Qualitative interviews reinforced these findings: feedback helped participants adopt a client-centered listening approach, while practice-alone participants remained solution-oriented. These results suggest LLM-based training systems can promote effective skill development, and combining simulated practice with structured feedback is critical for meaningful improvement.
Empirical studies in HCI, Interactive learning environments, LLM-based simulation
journalyear: 2026copyright: ccconference: Proceedings of the 2026 CHI Conference on Human Factors in Computing Systems; April 13–17, 2026; Barcelona, Spainbooktitle: Proceedings of the 2026 CHI Conference on Human Factors in Computing Systems (CHI ’26), April 13–17, 2026, Barcelona, Spaindoi: 10.1145/3772318.3791821isbn: 979-8-4007-2278-3/2026/04ccs: Human-centered computing Empirical studies in HCIccs: Human-centered computing Interactive systems and toolsccs: Human-centered computing Natural language interfacesccs: Computing methodologies Natural language processing
1. Introduction
@@
In 2023, 22.8% of U.S. adults (approximately 58.7 million people) experienced a mental illness (Substance Abuse and Mental Health Services Administration, 2024). Yet, access to effective mental health care is severely limited by shortages of qualified providers, from psychotherapists and counselors to social workers and peer supporters (Larson et al., 2017; Modi et al., 2022). While there is increasing interest in direct-to-patient AI systems with some promising results (Heinz et al., 2025), we expect the demand for human-delivered mental health support to continue to far exceed supply. The limited supply of effective therapy providers is due, at least in part, to the reliance on resource-intensive methods to train helping skills (Nelson-Jones, 2013) and evidence-based interventions (Cook et al., 2017; Frank et al., 2020) which require access to trainers who can simulate a client interaction (Kuehne et al., 2018; Ay-Bryson et al., 2023) and provide expert supervision (Wheeler and Richards, 2007; Watkins Jr and Milne, 2014), limiting training scale (Atkins et al., 2014; Kühne et al., 2020).
@@
AI systems have been increasingly applied to counselor training as a potential solution to these scaling challenges. Recent advances in large language models (LLMs) have enabled the simulation of patients seeking mental health support (Wang et al., 2024; Louie et al., 2024), offering rich opportunities for practice. The use of simulated patients is not new: in medical and nursing education, human role-plays and standardized patients are routinely used, and meta-analyses show they significantly improve skill acquisition and learner confidence (Shin et al., 2015). Mental health training has relied on a similar tradition of human role-plays to develop core helping skills. In parallel, AI feedback systems have progressed in automatically assessing counselor behaviors such as empathy, reflections, and active listening (Xiao et al., 2015; Sharma et al., 2020; Flemotomos et al., 2021; Shah et al., 2022), generating suggested responses (Sharma et al., 2023; Hsu et al., 2025; Chaszczewicz et al., 2024) and explanations (Rudolph et al., 2024; Chaszczewicz et al., 2024). These feedback systems target skills from client-centered approaches (Raskin and Rogers, 2005; Miller and Rollnick, 2012), empathy, reflections, questions, and active listening, which have been shown to strengthen common factors like therapeutic alliance, a powerful predictor of therapy outcomes across therapy modalities (Elliott et al., 2018; Wampold, 2015; Norcross and Lambert, 2018). However, most evaluations have largely positioned AI as a real-time co-pilot rather than a training tool (Sharma et al., 2023; Hsu et al., 2025), or studied pre-LLM training systems with limited practice realism and simpler, non-generative feedback mechanisms (Tanana et al., 2019).
Recent work has developed and tested systems integrating LLM-simulated patients for training counselors, to learn case conceptualization skills for cognitive behavioral therapy (CBT) (Wang et al., 2024) or motivational-interviewing techniques for substance-misuse scenarios (Steenstra et al., 2025). Their studies recruit student and expert counselors to evaluate the system output quality, usability, and desirability. While prior work highlights the promise of LLM-simulations for training counselors, to our knowledge, no study has evaluated whether these LLM-simulated training systems promote novice skill improvements. A key barrier to conducting skill development studies is first establishing the LLM-system’s output quality; even advanced prompt-based approaches (e.g., providing coding manuals like the Motivational Interviewing Treatment Integrity (MITI) in context (Steenstra et al., 2025)) require validation and refinement by domain-experts.